# Modeling Multi-Task Model Merging as Adaptive Projective Gradient Descent

### v1

@@ 1 Introduction
Fine-tuning on pre-trained foundational models to address downstream tasks has become an effective paradigm (Muqeeth et al., 2024). However, the independent deployment of multiple fine-tuned models increases storage costs. While traditional multi-task learning (MTL) can mitigate these issues, they typically require concurrent training across multiple task-specific datasets, which incurs significant training overhead and potential privacy risks (Wei et al., 2024b). Consequently, there is a growing interest in merging multiple expert models into a unified model without accessing their original data (Yang et al., 2024a; Huang et al., 2024). Model merging is performed directly at the parameter level and maintains only one final model during inference. In recent years, numerous pre-trained and fine-tuned checkpoints have been released on open-source communities like GitHub or Hugging Face, making it easy to obtain expert models from diverse domains. These rich model repositories underscore the value of model merging.
@@ 2 Related Work
Model merging. Model merging (Crisostomi et al., 2024; Wang et al., 2024) eliminates the need for raw training data or expensive computations. It operates directly at the parameter level and consolidates multiple models into a single final model for inference. Existing model merging methods are categorized into two paradigms: pre-merging and during-merging (Yang et al., 2024a). Pre-merging methods aim to create favorable conditions for merging, such as using linearized fine-tuning to achieve weight disentanglement (Ortiz-Jimenez et al., 2024; Tang et al., 2024b).
During-merging focuses on designing techniques to merge given models, and is further divided into data-free methods and test-time adaptation (TTA) methods. Data-free methods rely solely on pre-trained and fine-tuned models for weight merging (Choi et al., 2024). For example, Task Arithmetic (TA) (Ilharco et al., 2023) efficiently transfers capabilities by combining task vectors using arithmetic operations. Ties-Merging (Yadav et al., 2024) trims redundant parameters based on magnitudes, resolves parameter sign conflicts, and merges sparse models. Concrete Merging (Tang et al., 2023) introduces a meta-learning framework to learn a concrete mask that suppresses conflicting parameters.
TTA methods assume access to unlabeled test datasets, which can be considered a form of transductive learning. AdaMerging (Yang et al., 2024c) learns the merging coefficients by minimizing entropy as a surrogate loss on test data. Representation Surgery (Yang et al., 2024b) calibrates biases and aligns the representation of the merged model with that of independent models. In contrast, we design a data-free objective to resolve conflicts and optimize this objective within a shared subspace to retain shared knowledge.
Multi-task learning. Existing MTL research addresses the issue of negative transfer (Jiang et al., 2023) from two principal perspectives: architecture and optimization. From the architectural perspective, negative transfer is mitigated through strategies like modularization (Lu et al., 2024), sparsification (Sun et al., 2020), or soft sharing of the model backbone (Misra et al., 2016). From an optimization perspective, it is widely recognized that tasks sharing similar underlying structures can benefit from being trained together. Zamir et al. (2018) advocate for segregating unrelated tasks to minimize interference. Gradient alignment methods (Shi et al., 2022) emphasize maintaining consistency in gradient directions or signs to resolve conflicts, with techniques like PCGrad (Yu et al., 2020), which projects one task’s gradient onto the normal plane of another task’s conflicting gradient to reduce interference. Distinct from these traditional MTL approaches focused on addressing negative transfer, our method improves multi-task performance by aligning the merged model with each individual model through a detailed decomposition and reconstitution of the loss function.
@@ 3.2 Rethinking Model Merging for MTL
Resolving conflicts among tasks is a key challenge in model merging. Unlike MTL, which can mitigate conflicts during training with access to original data, model merging operates entirely in the parameter space. Existing methods mainly address conflicts by sparsely adjusting parameters, either by dropping conflicting parameters based on signs (Yadav et al., 2024) or importance scores (Du et al., 2024). Other methods promote orthogonality among task vectors, either by fine-tuning models in the tangent space (Ortiz-Jimenez et al., 2024) or directly optimizing task vectors (Xiong et al., 2024). While these methods alleviate conflicts, they inevitably discard task-specific information that contributes to conflicts, resulting in performance degradation. The fundamental requirement of model merging is that the merged model should perform comparably to task-specific models on their respective tasks—an aspect existing methods fail to consider. As shown in Fig. 1, increasing the task numbers leads to a continuous performance decline across methods. This is because more tasks result in increased negative transfer, causing the discard of valuable conflict-related task-specific knowledge. Therefore, we propose explicitly modeling the gap between the merged model $\boldsymbol{\theta}^{*}$ and individual models $\boldsymbol{\theta}_{i}$. This transforms conflict resolution into an optimization problem that can be solved using gradient descent.
@@ 4 Methodology
$\min_{\boldsymbol{\theta}^{*}}$ $\sum_{i=1}^{n}\|\boldsymbol{\theta}^{*}-\boldsymbol{\theta}_{i}\|_{Gap},$ (2)
@@
Then, we solve it via an adaptive projective gradient descent $\mathbf{g}=\mathrm{Proj}_{\perp S_{\text{share}}}\!\bigl{(}\nabla_{\Delta}\mathcal{L}(\Delta;\lambda_{i},\boldsymbol{\tau}_{i})\bigr{)}$ method.
@@ 4.1 A Data-Free ̵Objective
Existing methods (e.g., TA) can efficiently combine multiple task-specific models. However, these methods often exhibit notable performance gaps compared to single-task models. Considering the fundamental requirement that the merged model should perform comparably to its respective task-specific model for each task, we formally define the objective to resolve model merging as follows:
@@
where $\mathcal{L}_{j}(\boldsymbol{\theta})$ denotes the loss for task $j$ with model parameters $\boldsymbol{\theta}$. This objective requires that the merged model’s performance on each task closely aligns with the performance achieved when only the corresponding task vector $\boldsymbol{\tau}_{j}$ is integrated.
Multi-task conflicts often arise during model merging, as expert models encapsulate diverse and sometimes conflicting knowledge. Therefore, we introduce a modification vector $\Delta$ to each task vector, aiming to alleviate conflicts by optimizing $\Delta$. Previous work (Xiong et al., 2024) has shown that eliminating redundant components from task vectors can help reduce interference between tasks. In this context, $\Delta$ can be understood as the shared redundant portion of task vectors. However, directly optimizing Eq. 3 requires task-specific data to compute $\mathcal{L}_{j}$, which is unavailable as we only have access to model parameters. To overcome this limitation, we apply a Taylor expansion around the pre-trained model parameters $\boldsymbol{\theta}_{0}$:
@@ 6 Conclusion
Merging multiple expert models provides an efficient solution for multi-task learning without data. However, existing methods primarily focus on mitigating task conflicts, neglecting the fundamental requirement of model merging: ensuring comparable performance to task-specific models. In this paper, we frame model merging as a constrained optimization problem and address it using an adaptive projective gradient descent method, enabling data-free optimization of task vectors. Extensive experiments demonstrate the superior generalization and robustness of our approach, showcasing its effectiveness across diverse benchmarks.

### v2

@@ 1 Introduction
Fine-tuning on pre-trained foundational models to address downstream tasks has become an effective paradigm (Muqeeth et al., 2024). However, the independent deployment of multiple fine-tuned models increases storage costs. While traditional multi-task learning (MTL) can mitigate these issues, they typically require concurrent training across multiple task-specific datasets, which incurs significant training overhead and potential privacy risks (Wei et al., 2024). Consequently, there is a growing interest in merging multiple expert models into a unified model without accessing their original data (Yang et al., 2024a; Huang et al., 2024). Model merging is performed directly at the parameter level and maintains only one final model during inference. In recent years, numerous pre-trained and fine-tuned checkpoints have been released on open-source communities like GitHub or Hugging Face, making it easy to obtain expert models from diverse domains. These rich model repositories underscore the value of model merging.
@@ 2 Related Work
Model merging. Model merging (Crisostomi et al., 2024; Wang et al., 2024) eliminates the need for raw training data or expensive computations. It operates directly at the parameter level and consolidates multiple models into a single final model for inference. Existing model merging methods are categorized into two paradigms: pre-merging and during-merging (Yang et al., 2024a). Pre-merging methods aim to create favorable conditions for merging, such as using linearized fine-tuning to achieve weight disentanglement (Ortiz-Jimenez et al., 2023; Tang et al., 2024b).
During-merging focuses on designing techniques to merge given models, and is further divided into data-free methods and test-time adaptation (TTA) methods. Data-free methods rely solely on pre-trained and fine-tuned models for weight merging (Choi et al., 2024). For example, Ties-Merging (Yadav et al., 2024) trims redundant parameters based on magnitudes, resolves parameter sign conflicts, and merges sparse models. Concrete Merging (Tang et al., 2023) introduces a meta-learning framework to learn a concrete mask that suppresses conflicting parameters. Calculating the loss gap has been reflected in some studies: MetaGPT (Zhou et al., 2024) formally defines the loss difference, deriving a closed-form solution for $\lambda$. TATR (Anonymous, 2024) defines knowledge conflict between losses and expands it as the product of gradients and task vectors. We focus on optimizing task vectors via gradient descent while constraining it within a shared subspace to retain shared knowledge.
TTA methods assume access to unlabeled test datasets, which can be considered a form of transductive learning. AdaMerging (Yang et al., 2024c) learns the merging coefficients by minimizing entropy as a surrogate loss on test data. Representation Surgery (Yang et al., 2024b) calibrates biases and aligns the representation of the merged model with that of independent models. In contrast, we design a data-free objective to resolve conflicts.
Multi-task learning. Existing MTL research addresses the issue of negative transfer (Jiang et al., 2023) from two principal perspectives: architecture and optimization. From the architectural perspective, negative transfer is mitigated through strategies like modularization (Lu et al., 2024), sparsification (Sun et al., 2020), or soft sharing of the backbone. From an optimization perspective, it is widely recognized that tasks sharing similar underlying structures can benefit from being trained together. Gradient alignment methods (Yu et al., 2020; Shi et al., 2022) emphasize maintaining consistency in gradient directions or signs to resolve conflicts, which projects one task’s gradient onto the normal plane of another task’s gradient to reduce forgetting (Saha et al., 2021). Our approach enhances multi-task performance by aligning the merged model with each individual model and utilizing adaptive merging coefficients.
@@ 3.2 Rethinking Model Merging for MTL
Resolving conflicts among tasks is a key challenge in model merging. Unlike MTL, which can mitigate conflicts during training with access to original data, model merging operates entirely in the parameter space. Existing methods mainly address conflicts by sparsely adjusting parameters, either by dropping conflicting parameters based on signs (Yadav et al., 2024) or importance scores (Du et al., 2024). Other methods promote orthogonality among task vectors, either by fine-tuning models in the tangent space (Ortiz-Jimenez et al., 2023) or directly optimizing task vectors (Xiong et al., 2024). While these methods alleviate conflicts, they inevitably discard task-specific information that contributes to conflicts, resulting in performance degradation. The fundamental requirement of model merging is that the merged model should perform comparably to task-specific models on their respective tasks—an aspect existing methods fail to consider. As shown in Fig. 1, increasing the task numbers leads to a continuous performance decline across methods. This is because more tasks result in increased negative transfer, causing the discard of valuable conflict-related task-specific knowledge. Therefore, we propose explicitly modeling the gap between the merged model $\boldsymbol{\theta}^{*}$ and individual models $\boldsymbol{\theta}_{i}$. This transforms conflict resolution into an optimization problem that can be solved using gradient descent.
@@ 4 Methodology
$\min_{\boldsymbol{\theta}^{*}}$ $\mathcal{L}(\Delta;\lambda_{i},\boldsymbol{\tau}_{i}):=\sum_{i=1}^{n}\|\boldsymbol{\theta}^{*}-\boldsymbol{\theta}_{i}\|_{\text{Gap}},$ (2)
@@
Then, we solve it via adaptive projective gradient descent:
$\Delta=\Delta-\mathbf{g},\text{where}\,\,\mathbf{g}=\mathrm{Proj}_{\perp S_{\text{share}}}\!\bigl{(}\nabla_{\Delta}\mathcal{L}(\Delta;\lambda_{i},\boldsymbol{\tau}_{i})\bigr{)}.$
It uses adaptive $\lambda_{i}$ for different tasks, projects the gradient orthogonal to $S_{\text{share}}$ to satisfy the constraint, and optimizes the modification of $\boldsymbol{\theta}^{*}$ to minimize the loss.
@@ 4.1 A Data-Free ̵Objective
Considering the fundamental requirement that the merged model should perform comparably to its respective task-specific model for each task, we follow Zhou et al. (2024) to define the objective for resolving model merging as:
@@
where $\mathcal{L}_{j}(\boldsymbol{\theta})$ denotes the loss for task $j$ with model parameters $\boldsymbol{\theta}$. This objective requires that the merged model’s performance on each task closely matches the performance achieved using only the corresponding task vector $\boldsymbol{\tau}_{j}$.
Multi-task conflicts often arise during model merging, as expert models encapsulate diverse and sometimes conflicting knowledge. Therefore, we introduce a modification vector $\Delta$ to each task vector, aiming to alleviate conflicts by optimizing $\Delta$. Previous work (Xiong et al., 2024) has shown that eliminating redundant components from task vectors can help reduce interference between tasks. In this context, $\Delta$ can be understood as the shared redundant portion of task vectors. However, directly optimizing Eq. 3 requires task-specific data to compute $\mathcal{L}_{j}$, which is unavailable as we only have access to model parameters. To overcome this limitation, we apply a Taylor expansion around the pre-trained model parameters $\boldsymbol{\theta}_{0}$ (Ortiz-Jimenez et al., 2023):
@@ 6 Conclusion
Existing merging methods often prioritize mitigating task conflicts, neglecting a critical requirement of model merging: achieving performance comparable to task-specific models. In this paper, we rethink model merging from a multi-task learning perspective, treating it as a constrained optimization problem. We introduce an adaptive projective gradient descent method that optimizes a data-free objective within a shared subspace and includes adaptive merging coefficients. Extensive experiments validate the superior generalization and robustness of our approach, highlighting its effectiveness across various benchmarks.
@@ Appendix D Experiment Results
More task numbers.
Fig. 1 illustrates the robustness of our approach when handling a larger number of tasks. Following Wang et al. (2024), we evaluate its performance as more tasks are merged. In addition to the previously used 8 tasks, the 14-task scenario incorporates CIFAR100, STL10, Flowers102, OxfordIIITPet, PCAM, and FER2013. The 20-task scenario further adds six tasks: EMNIST, CIFAR10, Food101, FashionMNIST, RenderedSST2, and KMNIST. Our approach exhibits increasingly significant performance advantages as the number of tasks grows, demonstrating its effectiveness in mitigating negative transfer through gradient descent while preserving task-specific knowledge.
Table 11: Average accuracy (%) when merging models across a larger number of tasks.
Method ViT-B/32 ViT-L/14
8 tasks 14 tasks 20 tasks 8 tasks 14 tasks 20 tasks
Pre-trained 48.4 57.3 56.1 64.4 68.0 65.1
Weight averaging 66.5 64.4 61.1 79.4 76.6 71.5
Task arithmetic 70.8 65.4 60.6 84.8 79.3 74.0
TIES 75.1 68.0 63.4 86.9 79.5 75.7
Consensus TA 75.0 70.4 65.4 86.2 82.2 78.9
Data-Free Consensus TIES 74.8 67.7 63.2 86.9 81.5 76.8
doGe TA (Ours) 80.7 (+5.7%) 77.9 (+7.5%) 72.5 (+7.1%) 88.8 (+2.6%) 87.1 (+4.9%) 81.0 (+2.1%)